OpenStudy (anonymous):

Solve the system using substitution. 5x = -14 + y 3y = 20 + 4x

jimthompson5910 (jim_thompson5910):

5x = -14 + y 5x +14 = y y = 5x + 14 ---------------- 3y = 20 + 4x 3(5x+14) = 20 + 4x 15x+42 = 20 + 4x I'll let you finish

OpenStudy (anonymous):

y= 5x +14 insert at eq2 3(5x+14)= 20+4x 15x+42=20 +4x x=2 y=5(2)+14 y=24

OpenStudy (ryan123345):

So @jim_thompson5910 just went all the way to 15x + 42 = 20 + 4x. Just combine like terms and isolate the variable :)

OpenStudy (anonymous):

doesnt x= -2 and y = 4.... i may be wrong?

jimthompson5910 (jim_thompson5910):

yes, x = -2 and y = 4

OpenStudy (anonymous):

x=-2 y=4